# Efficient Nearest Neighbor based Uncertainty Estimation for Natural Language Processing Tasks

### v1

@@ Appendix A Training Settings for Density Estimator in Density Softmax
Appendix B Details of Evaluation Metrics
@@
MCE, as detailed by Naeini et al. (2015) measures the maximum difference between the model’s accuracy and the confidence across variouts confidence levels. MCE is defined as:
@@
Appendix C Each Result of Product Quantization, Clustering, and Dimension Reduction
C.1 Product Quantization
@@
We evaluated uncertainty estimation performance and inference time when the number of clusters in the codebook was fixed at 32, and the number of subvectors was changed to .
@@
Table 11 shows the uncertainty estimation performance and inference time results in different . In ECE and E-AURC, there are almost no degradation in uncertainty estimation performance due to PQ. On the other hand, in MCE in ID setting, the uncertainty estimation performance consistently degrades. Furthermore, compared to NN-UE among different , the larger , the better the uncertainty estimation performance tends to improve, but the inference time increases.
@@
C.2 Clustering
In this study, we evaluate the uncertainty estimation performance and inference speed when the number of clusters is fixed and number of cluster centroids to search changes.
Table 12 shows the performance of uncertainty estimation when changing in ID and OOD settings using OntoNotes 5.0. In ECE, scores are slightly reduced for ID, but only slightly worse for OOD; MCE also shows degradation for ID but little for OOD, and even improves when ; E-AURC shows almost no change in scores when is changed for both ID and OOD. In terms of inference time, the larger , the longer it takes.
@@ C.2 Clustering
C.3 Dimension Reduction
As shown in Table 13, the uncertainty estimation performance depends on the number of target dimension, and the performance degrades when or . On the other hand, the performance in is almost the same as default NN-UE. This is because while dimension reduction mitigates the problem of high dimensionality, the failure to extract adjacent examples by NN search when applying PCA is less likely compared to PQ and clustering if the target dimension in dimension reduction is not reduced too much.
@@ C.3 Dimension Reduction
Appendix D Distance Recomputation for NN-UE
When using efficient NN search techniques in Section 7.2, we use approximate distances to compute Eq. 5. Although we can get raw vectors by using the example indices obtained from approximate nearest neighbor search and compute accurate distance, in NN-LM this has been shown to lead to performance gains and latency degradation He et al. (2021a). We measure the uncertainty estimation performance and inference speed when PQ, clustering, and dimension reduction are applied simultaneously and re-computing accurate distances, reported in Table 14. These results show that the uncertainty estimation performance does not improve except for MCE in the ID setting, and the latency is about 5-7x slower when reading raw vectors from the datastore and re-computing distances.
@@ Appendix D Distance Recomputation for NN-UE
Appendix E Licenses of Datasets, Tools and Models
@@
IMDb movie dataset can be used for research purpose as described in https://developer.imdb.com/non-commercial-datasets/. Yelp Polarity dataset can be used for academic purpose as described in https://s3-media0.fl.yelpcdn.com/assets/srv0/engineering_pages/f64cb2d3efcc/assets/vendor/Dataset_User_Agreement.pdf. MNLI dataset is licensed for research purpose as described in Williams et al. (2018). SNLI dataset can be used for research purpose as described in https://nlp.stanford.edu/projects/snli/. OntoNotes 5.0 dataset can be used for research purpose as described in https://catalog.ldc.upenn.edu/LDC2013T19.

### v2

@@ Appendix B Training Settings for Density Estimator in Density Softmax
Appendix C Details of Baselines
Softmax Response (SR)
Temperature Scaling (TS)
is a calibration technique by which the logits are divided by a temperature parameter before applying the softmax function Guo et al. (2017). We optimized by L-BFGS on validation set loss.
Label Smoothing (LS)
MC Dropout
is an UE technique by times stochastic inferences with activating dropout Gal and Ghahramani (2016). In our experiments, we set for all evaluations, and the dropout rate is 0.1.
Spectral-Normalized Gaussian Process (SNGP)
uses spectral normalization of the weights for distance-preserving representation and Gaussian Processes in the output layer for estimating uncertainty Liu et al. (2020).
Posterior Networks (PN)
Mahalanobis Distance with Spectral-Normalized Network (MDSN)
is a Mahalanobis distance based UE method that benefits from by spectral normalization of the weights Vazhentsev et al. (2022), similar to SNGP.
E-NER
applies EDL framework for NER by introducing uncertainty-guided loss terms Zhang et al. (2023).
Appendix D Details of Evaluation Metrics
@@
MCE, as detailed by Naeini et al. (2015) measures the maximum difference between the model’s accuracy and the confidence across variousb confidence levels. MCE is defined as:
@@
Appendix E Additional Results on the Brier score
The Brier score is a widely used metric in UE community for evaluating the probabilistic predictions. The metric measures the mean squared difference between the predicted probability assigned to the predicted label and the actual outcome. This evaluation serves as a holistic assessment of model performance, reflecting both fit and calibration, in the following formula:
| Brier score | (10) |
where is the predicted probability assigned to the prediction, and is the actual outcome. Table 11 shows the results on the Brier score. These results indicate NN-UE improves calibration performance more prominently than other methods while maintaining prediction performance.
| Methods | SA | NLI | ||
| IMDb | Yelp | MNLI | SNLI | |
| SR | 5.000.27 | 5.830.98 | 9.500.40 | 11.020.41 |
| TS | 5.090.42 | 6.671.36 | 8.310.25 | 9.600.21 |
| LS | 4.640.23 | 5.160.92 | 8.730.23 | 10.180.17 |
| MC Dropout | 4.880.21 | 5.450.55 | 9.330.36 | 11.000.28 |
| SNGP | 4.780.15 | 5.990.39 | 12.255.38 | 13.454.57 |
| PN | 10.310.28 | 11.160.22 | 20.760.32 | 21.110.42 |
| Density Softmax | 4.820.18 | 6.050.38 | 9.600.34 | 11.280.41 |
| DAC | 4.440.33 | 5.440.71 | 8.210.25 | 9.550.35 |
| kNN-UE (w/o label) | 4.370.16 | 5.100.12 | 8.150.15 | 9.520.32 |
| kNN-UE | 4.210.14 | 5.020.42 | 8.070.18 | 9.440.28 |
Appendix F The impact of NN-UE with label term in NER on E-AURC
NER tasks are often in label imbalanced settings, where the "O" label is typically much more than other entity-related labels. Additionally, in the OntoNotes 5.0 dataset, the number of labels is 37, as shown in Table 9, which is significantly higher than in SA and NLI tasks. As a result, compared to SA and NLI, neighbor labels will contain much more different labels from the predicted label. The presence of many other labels in the neighbors that are different from the predicted label can lead to excessively low confidence in NN-UE using the label term, even though the prediction is correct because in the label term becomes lower in NER. The impact of that bias for calibration errors, such as ECE and MCE, will be limited. However, low confidence in accurate prediction reduces the coverage much in E-AURC, leading to a degradation in E-AURC.
For example, assume that in a test data set of 6 examples for 3 classes classification, the predictions for the first 3 examples are incorrect and the latter 3 examples are correct. In case A, we assume that prediction confidences are [[0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.02, 0.02, 0.96], [0.01, 0.01, 0.98]]. In case B, we assume that prediction confidences are [[0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.25, 0.25, 0.5], [0.275, 0.275, 0.45], [0.02, 0.02, 0.96], [0.01, 0.01, 0.98]]. In these settings, the ECE and E-AURC for each case are shown in Table 12. These scores indicate that E-AURC is strongly penalized when the confidence in a correct prediction is lower than the confidence in an incorrect prediction.
| Case | ECE (↓) | E-AURC (↓) |
| A | 17.67 | 18.80 |
| B | 16.83 | 121.57 |
Appendix G Inference Time Full Results
We show the inference time full results on out-of-domain test sets in Table 13.
| Methods | SNLI | OntoNotes 5.0 |
| SR | 21.590.76 | 5.750.27 |
| TS | 21.640.07 | 5.790.17 |
| LS | 21.700.07 | 5.800.19 |
| MC Dropout | 396.861.10 | 101.980.83 |
| SNGP | 24.590.08 | - |
| PN | 23.260.05 | - |
| MDSN | 23.390.85 | - |
| E-NER | - | 5.780.61 |
| Density Softmax | 22.020.05 | 6.020.07 |
| DAC | 2346.6236.06 | 326.001.41 |
| NN-UE (w/o label) | 23.020.04 | 10.360.21 |
| NN-UE | 23.070.05 | 10.480.12 |
Appendix H Inference Time Results When Changing -
To estimate whether the inference time changes significantly when changing Top- in kNN search, we investigated the inference time when changing on the IMDb test set. Table 14 shows that the inference time remains almost the same when changing in the range of 8 to 128.
| Methods | Inference time [s] |
| SR | 121.560.12 |
| NN-UE (=8) | 128.980.11 |
| NN-UE (=16) | 128.970.13 |
| NN-UE (=32) | 128.540.16 |
| NN-UE (=64) | 129.160.16 |
| NN-UE (=128) | 128.390.20 |
@@ Appendix H Inference Time Results When Changing -
Appendix I Each Result of Product Quantization, Clustering, and Dimension Reduction
I.1 Product Quantization
(PQ) Jégou et al. (2011) is a data compression technique based on vector quantization. In PQ, a -dimensional representation is divided into subvectors and quantized by performing -means clustering on the vectors in each subspace. Vector quantization can significantly reduce the amount of memory occupied by vectors.131313For example, raw datastore in kNN-UE is 636MB on OntoNotes 5.0 , but PQ reduces it to 10MB. In addition, by calculating the distance between compressed PQ codes, we can efficiently calculate the estimated value of the original Euclidean distance.
@@ I.1 Product Quantization
We evaluated UE performance and inference time when the number of clusters in the codebook was fixed at 32, and the number of subvectors was changed to (In Table 7 and 8, PQ was performed with ).
@@
Table 15 shows the UE performance and inference time results in different . In ECE and E-AURC, there are almost no degradation in UE performance due to PQ. On the other hand, in MCE in ID setting, the UE performance consistently degrades. Furthermore, compared to NN-UE among different , the larger , the better the UE performance tends to improve, but the inference time increases.
The larger is, the more time is required for inference but the UE performance improves. We assumed that these results are derived from the decrease in quantization error over the vector with PQ with larger because each subvector is divided into smaller subspaces and the quantization is performed for each subspace. On the other hand, an increase in requires additional distance computations etc., then more inference time.
@@
I.2 Clustering
The original NN-LM uses an inverted file index (IVF) technique that speeds up the search by dividing the representation into clusters by -means and searching for neighbors based on centroids. In this study, we evaluate the UE performance and inference speed when the number of clusters . In this study, we evaluate the UE performance and inference speed when the number of clusters and applying PQ with are fixed and the number of cluster centroids to search changes (In Table 7 and 8, IVF was performed with ).
Table 16 shows the performance of UE when changing in ID and OOD settings using OntoNotes 5.0. In ECE, scores are slightly reduced for ID, but only slightly worse for OOD; MCE also shows degradation for ID but little for OOD, and even improves when ; E-AURC shows almost no change in scores when is changed for both ID and OOD. In terms of inference time, the larger , the longer it takes. We derive the improvement in MCE when increasing in ID setting from the fact that more clusters are targeted, making it possible to cover ground-truth nearest neighbor examples. On the other hand, the tendency of slight decrease when increasing in OOD setting may comes from the reliability of the vector, similar to the discussion in Section 6.3.
In addition, Taken together with the results in Table 7 in Section 7.3, we can see that the degradation of the UE performance can be mitigated with improvement latency when applying PQ and IVF with lower , compared to applying PQ, IVF and PCA simultaneously.
@@ I.2 Clustering
I.3 Dimension Reduction
In general, Transformer-based models such as PLM have high-dimensional token representations. In high-dimensional spaces, nearest neighbor search often suffer from the curse of dimensionality. To reduce this problem, we apply dimension reduction to NN-UE similar to He et al. (2021a). In this study, we use Principal Component Analysis (PCA) as a dimension reduction algorithm to reduce the dimension of the datastore representations and the query representation (In Table 7 and 8, PCA was performed with ). As shown in Table 17, the UE performance depends on the number of target dimensions, and the performance degrades when or . On the other hand, the performance in is almost the same as default NN-UE. This suggests that excessive dimension reduction in distance computation to extract nearest examples by NN search could have a negative impact on the UE performance.
@@ I.3 Dimension Reduction
Appendix J Distance Recomputation for NN-UE
When using efficient NN search techniques in Section 7.3, we use approximate distances to compute Eq. 4. Although we can get raw vectors by using the example indices obtained from approximate nearest neighbor search and compute accurate distance, in NN-LM this has been shown to lead to performance gains and latency degradation He et al. (2021a). We measure the UE performance and inference speed when PQ, clustering, and dimension reduction are applied simultaneously and re-computing accurate distances, reported in Table 18. These results show that the UE performance does not improve except for MCE in the ID setting, and the latency is about 5-7x slower when reading raw vectors from the datastore and re-computing distances. Moreover, these results suggest that exact distance computation for examples that are not actually nearest neighbors are not very effective in NN-UE.
@@
Appendix K Licenses of Datasets, Tools and Models
@@
The IMDb movie dataset can be used for research purposes as described in https://developer.imdb.com/non-commercial-datasets/. Yelp dataset can be used for academic purposes as described in https://s3-media0.fl.yelpcdn.com/assets/srv0/engineering_pages/f64cb2d3efcc/assets/vendor/Dataset_User_Agreement.pdf. The MNLI dataset is licensed for research purposes as described in Williams et al. (2018). The SNLI dataset can be used for research purposes as described in https://nlp.stanford.edu/projects/snli/. OntoNotes 5.0 dataset can be used for research purposes as described in https://catalog.ldc.upenn.edu/LDC2013T19.